the width of a rectangular swimming pool is 12 meters less than the length and the perimeter is 72 meters find its dimensions
Answer by ikleyn(52776)

.
W + L + W + L = 72   perimeter


(L-12) + L + (L-12) + L = 72


4L - 24 = 72


etc.